OpenStudy (anonymous):

Find exact value in radians: inverse tan of [tan(2pi/5)]?

OpenStudy (anonymous):

\[\tan^{-1} (\tan 2\pi/5)\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

wait i am solving.......

OpenStudy (anonymous):

is 72 degrees the answer?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

no has to be in radians

OpenStudy (anonymous):

i dont have the answer though

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

its so simple the answer is 2pi/5

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

im trying to come up with that, how do you end up with 2pi/5?

OpenStudy (anonymous):

actually \[\tan^{-1} (\tan (\theta))=\theta \] provided \[\theta lies \between -\pi/2 \to \pi/2\]

OpenStudy (anonymous):

provided theta lies between -pi/2 to pi/2

OpenStudy (anonymous):

oh wow haha i cant believe i forgot that property

OpenStudy (anonymous):

are you making fun of me or reminding yourself?

OpenStudy (anonymous):

no not at all im laughing at myself cause youre right it was so simple!

OpenStudy (anonymous):

ok any other help required?

OpenStudy (anonymous):

Nope that will be all, thanks!